Shift report 20010930 16:00-24:00

From: Apache (apache@pii3.brahms.bnl.gov)
Date: Sun Sep 30 2001 - 00:03:42 EDT

  • Next message: Apache: "Shift report 20010930 00:00-08:00"

    ------------------------------------------------------------
    Date: Sun Sep 30 00:03:42 2001, Shift: 16:00-24:00
    Supervisor: sjs, with: 
    Log book#9 pages 65 - 67
    
    Shift Summary:
    ------------------------------------------------------------
    1. Usable beam from 17:10-19:55.  Then "Quench lock"+
       "control problems" for last half of shift.
    2. Runs 5045 + 5046 finish 90deg "4A" setting.   
       D5 set to B polarity after run 5046.  
       FV modified DAQ to read new calibration fibers in
       front of tpm2.
       Note: 5037 of previous shift has mislabelled   
       D5 polarity in DAQ DB:  The run had D5 in "A" polarity,
       not "B", as entered in DB.
    3. No new problems found. 
    ------------------------------------------------------------
    



    This archive was generated by hypermail 2b30 : Sun Sep 30 2001 - 00:04:30 EDT